OpenStudy (anonymous):

Solve for n: (4^n)^3 (2^n + 2) = (42)^n – 1. A. –2 B. –1 C. 0 D. 2

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

for clarification's sake \[\huge (4^n)^3 (2^{n+2}) = 42^{n-1}\] yes?

OpenStudy (anonymous):

\[(4^{n})^{3} (2^{n+2}) = (4^{2}) ^{n-1}\]

OpenStudy (anonymous):

that makes it a lot easier

OpenStudy (souvik):

you can use trail and error method..

OpenStudy (anonymous):

first, simplify the (4^n )^3 so you have 4^(3n) agree?

OpenStudy (anonymous):

(trial and error works too since this is multiple choice btw)

OpenStudy (anonymous):

after simplifying 4^3n, how do I continue?

OpenStudy (souvik):

4^(3n)=2^6n....agree?

OpenStudy (anonymous):

^that

OpenStudy (anonymous):

put everything in base of 2

OpenStudy (anonymous):

why 2^6n? I dont understand

OpenStudy (souvik):

\[4^{3n}=(2^{2})^{3n}=2^{6n}\]

OpenStudy (souvik):

\[(a ^{x})^{y}=a ^{xy}\]

OpenStudy (anonymous):

Ok I understand now!

OpenStudy (anonymous):

\[2^{6n} x 2^{n+2} = (4^{2})^{n-1}\]

OpenStudy (anonymous):

How do I continue?

OpenStudy (souvik):

now...\[a ^{x}*a^{y}=a ^{x+y}\]

OpenStudy (anonymous):

I dont understand! :(

OpenStudy (souvik):

okay...tell me a*a*a*a=?

OpenStudy (souvik):

no..its a*a*a*a=a^4

OpenStudy (anonymous):

6n x 2n + 2 = 2n-1

OpenStudy (anonymous):

is this correct?

OpenStudy (souvik):

what you said is ....a+a+a+a=4a

OpenStudy (anonymous):

Yes I understand the mistake I made

OpenStudy (souvik):

okay now \[2^{6n}*2^{n+2}=?\]

OpenStudy (anonymous):

6n + n+2

OpenStudy (souvik):

well done...its 2^(6n+n+2)=2^(7n+2)

OpenStudy (souvik):

now right hand side... \[4^{2}^(n-1) = ?\]

OpenStudy (souvik):

it is like...\[(a ^{x})^{y}=a ^{xy}\]

OpenStudy (anonymous):

4^2n-2 ?

OpenStudy (souvik):

good...now \[(2^{2})^{2n-2}=?\]

OpenStudy (anonymous):

2^4n+4 ?

OpenStudy (souvik):

there is minus sign...so 2^(4n-4) now...7n+2=4n-4.....as they all in equal base...

OpenStudy (anonymous):

is -2 the answer?

OpenStudy (souvik):

yep..correct !

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (souvik):

you are welcome...